OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. P(6, 2), Q(8, -4)

OpenStudy (anonymous):

y=mx+c m is gradient so m=(y1-y2)/(x1-x2) m=(2-(-4))/(6-8) m=-3 now, we put one point into equation, so lets take P, 2=-3(6)+c 2=-18+c 20=c so the equation is, y=-3x+20